OpenStudy (anonymous):

Hello, im new to this and am a yr 9 highschool student, at the moment we are studying Linear equations in algebra and im completely confused Here are some examples of problems i dont understand: Make the letter indicated in brackets the subject of the formula: 3t2(to the power of 2) -q= 12. I really dont get it, if someone were to answer this question i would be really greatful and ask more questions ;) !!!

OpenStudy (lgbasallote):

are there any preceding statements?

OpenStudy (anonymous):

Not really, no.

OpenStudy (lgbasallote):

hmm it says to make the letter in the bracekts the subject of the formula...does it look like this \[ [3t^2]^2 - q = 12\] exactly like that?? i mean there are brackets surrounding \(3t^2\)

OpenStudy (anonymous):

I wrote to the power of two because i didnt know how to type it, i just meant 3t2-q= 12 but with the 2 being the power, if that makes sense ;) !

OpenStudy (lgbasallote):

no no...the sense of my question was if there were brackets surrounding 3t^2 because the question says make the VARIABLE IN THE BRACKET the subject of formula...you getting my logic?

OpenStudy (anonymous):

My bad, the (q) was the letter in the brackets e.g. 3t2-q=12 Then it said "Make the letter indicated in brackets the subject of the formula, which is (q). -_-

OpenStudy (lgbasallote):

therefore we make q the subject of the formula

OpenStudy (anonymous):

Precisely.

OpenStudy (lgbasallote):

first get rid of \(3t^2\) how do you think that will be possible??

OpenStudy (anonymous):

No clue.... possibly though cancel out???

OpenStudy (anonymous):

Im guessing you put the q first as the subject for example: q=

OpenStudy (lgbasallote):

yes you are right

OpenStudy (lgbasallote):

to do that..we need to make 3t^2 be zero

OpenStudy (anonymous):

Ok, how?

OpenStudy (lgbasallote):

(3t^2)^2 - q ^what is the operation there?

OpenStudy (lgbasallote):

i mean between (3t^2)^2 and q..what is the operation??

OpenStudy (anonymous):

3t^2-q=12 was the sum....

OpenStudy (anonymous):

Then make (q) the subject...

OpenStudy (anonymous):

After that i didnt really get it..

OpenStudy (lgbasallote):

just focus on (3t^2)^2 - q...what is the operation? addition? subtraction? multiplication? division? others?

OpenStudy (anonymous):

DUDE DUDE DUDE the sum wasnt (3t^2)^2-q it was 3t^2-q=12 then make (q) the subject...

OpenStudy (lgbasallote):

oh woops...

OpenStudy (anonymous):

I dont think i was very clear, sorry...

OpenStudy (lgbasallote):

okay then... 3t^2 - q = 12..what is the operation?

OpenStudy (anonymous):

- ??

OpenStudy (lgbasallote):

what is the operation in 3t^2 - q is it addition? subtraction? multiplication? division? others?

OpenStudy (anonymous):

subtraction.... ?

OpenStudy (lgbasallote):

good! and what is the opposite of subtraction?

OpenStudy (anonymous):

addition?

OpenStudy (lgbasallote):

yes yes!! now i want you to look at this...you know that you need to make it "q =" right? let's say you have 2 - q = 1 how do you get rid of 2?

OpenStudy (anonymous):

q=2-1.... cancel out??? Dont really know....

OpenStudy (lgbasallote):

that's right actually...but just focus on 2- q...how do you get rid of 2?

OpenStudy (lgbasallote):

how can you make 2- q become 0 - q?

OpenStudy (anonymous):

Can you go q-2 then subtract resulting in no 2????

OpenStudy (lgbasallote):

do you mean 2 - q - 2?

OpenStudy (anonymous):

No, just q - 2....

OpenStudy (lgbasallote):

hmm no... how man you make 2 become 0?

OpenStudy (lgbasallote):

how can*

OpenStudy (anonymous):

I dont know... Explain?

OpenStudy (lgbasallote):

1 - 1 = 0 2- 2 = 0 3 - 3 = 0 remember?

OpenStudy (anonymous):

Ok, so to get 2-q you go 2-2= 0-q, so does that mean q is 2...

OpenStudy (lgbasallote):

no no..you're not looking for q...you were just finding a way to make 2 become 0

OpenStudy (lgbasallote):

look at 3t^2 - q how can you make 3t^2 0?

OpenStudy (anonymous):

2-2=0

OpenStudy (lgbasallote):

yes that's how you make 2 become 0..how do you make 3t^2 zero then?

OpenStudy (anonymous):

Ummmmm..... ????

OpenStudy (anonymous):

No clue...

OpenStudy (anonymous):

This is what i didnt really understand..

OpenStudy (lgbasallote):

you make 2 become 0 by subtracting 2 you make 1 become 0 by subtracting 1

OpenStudy (anonymous):

So i make 3t^2 become 0 by???

OpenStudy (lgbasallote):

think about it... 1- 1 = 0 2 - 2 = 0 x - x = 0 3t - 3t = 0

OpenStudy (anonymous):

3t^2-3t^2= 0!!!

OpenStudy (lgbasallote):

YES!!!

OpenStudy (lgbasallote):

now 3t^2 - q - 3t^2 = 12 BUT since you subtract 3t^2 from the left side..you also subtract 3t^2 to the right side for it to be equal make sense?

OpenStudy (lgbasallote):

therefore -q = 12 - 3t^2 you get that??

OpenStudy (anonymous):

I THINK SO!!!!

OpenStudy (lgbasallote):

now you need q = you have -q so you multiply everything by -1 to make q positive..make sense? q = -12 + 3t^2

OpenStudy (anonymous):

YEAH, I THINK SO, IT SOUNDS LIKE SOMETHING MY TEACHER WAS SAYING!!!

OpenStudy (anonymous):

so, QUICK RE-CAP...

OpenStudy (lgbasallote):

3t^2 - q = 12 3t^2 - q - 3t^2 = 12 - 3t^2 -q = 12 - 3t^2 q = -12 + 3t^2

OpenStudy (anonymous):

3T^2-Q=12 = 3T^2-Q-3T^2=12 = -Q=12-3T^2 = Q=-12 + 3T^2

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

Soz, my internet dropped out, it does that here... you live in australia?

OpenStudy (lgbasallote):

no

OpenStudy (anonymous):

Where is Malin anyways?

OpenStudy (anonymous):

Well, thanks for helping me with that sum...

OpenStudy (anonymous):

Bye, if its ok i may talk to you tommorrow about more sums!